is ball on Moon, where the acceleration due to gravity is 1/6(g -9.8) and air resistance is negligible, is hit directly upward and returns to the same level 85 s later, (a) How high above its origjinal point did the ball go? (b) How fast it moving just a after it was hit? 7. Two train cars are on a straight, horizontal track. The car A starts at rest and is put in motion with a constant acceleration of 2.00 m/s2. This car A moves toward the car B that is 25.0 m away. The car B is moving in same direction as car A with an initial velocity of 4.5 m/s and is traveling with a constant acceleration of -1.50 m/s2 a) Where will the cars collide? b) How long will it take for the cars to collide?
Answer #1

6. a) Hmax = 0.5*(g/6)*(t/2)2

Hmax = 0.5*1.63*4.252
Hmax = 14.7 m

b) v = (g/6)*(t/2)
v= 6.94 m/s